MFT(1) MFT(1) NAME mft−translateMetafontcodetoTeXcodeforprettyprinting SYNOPSIS mftmffilename[.mf][[−c]or[changefilename[.ch]]][[−s]or[stylefilename[.mft]]] DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualWeb2C:ATeXimplementation. ThemftprogramcreatesaTEXfilefromaMETAFONTprogram.Ittakesappropriatecareof typographicdetailslikepagelayoutandtheuseofindentation,italics,boldface,etc.,asillustrated inthebookComputerModernTypefaces.SpecialconventionsinMETAFONTcommentsallowyou tocontrolthingsthatwouldnototherwisecomeoutright;section1oftheMFTsourceprogram intheMETAFONTwarereportexplainstheserules. Thecommandlinehasonerequiredfilenameandtwooptionalfilenames.Therequiredoneisa METAFONTsourcefile;thereisalsoanoptionalchangefile(whichworksjustasthechangefilesto tangle(1)andweave(1)do)andanoptionalstylefile(whichisprependedtoeverything). Afilenamethatdoesn’tcontainadotisalwaysgivenanextension,either.mf(METAFONT)or.ch (change)or.mft(style).Ifnostylefileisspecified,thestylefileplain.mftisautomaticallyused. Thechangefilenameisnotsearchedforusinganypaths.The.mffileissearchedforusingthe MFINPUTSenvironmentvariableifyouhavesetit,orelsethesystemdefault.Thestylefileis searchedforusingtheTEXINPUTSenvironmentvariable.Seetex(1)forthedetailsofthe searching. TheoutputTEXfilenameisformedbyusing.texinplaceoftheextensionofmffilename. OPTIONS Specifycommand-lineoptions−ctosuppressthereadingofachangefile,and−stosuppressthe readingofastylefile. FILES mftmac.tex TEXmacrosusedbymftoutput. plain.mft Defaultstylefile. cmbase.mft StylefileforComputerModern. SEEALSO mf(1),weave(1). DonaldE.Knuth,ComputerModernTypefaces(VolumeEofComputersandTypesetting),Addison-Wesley,1986,ISBN0-201-13446-2. DonaldE.Knuthetal.,METAFONTware. AUTHORS DonaldE.Knuthwrotetheprogram,andheportedittoUnixwiththehelpofPierreMacKay andtheUnixportofweavebyHowardTrickeyandPavelCurtis.Theprogramispublishedin theMETAFONTwaretechnicalreport,availablefromtheTEXUsersGroup. Web2C7.32x 14December1993 1